# Dracula’s Chivito: discovery of a large edge-on protoplanetary disk with Pan-STARRS

### v1

@@ 4 Distance
Since DraChi is not associated with any known star-forming region the distance cannot be precisely determined, similar to GoHam which has estimates ranging from 250 to 500 pc. However, new stars form in interstellar clouds so we can look for the cloud associated with DraChi based on extinctions. We used Gaia DR3 extinctions estimates for stars located within 3 distance from DraChi. The Gaia extinctions were converted to Av using the dustapprox tool555https://github.com/mfouesneau/dustapprox. The result is shown in Figure 3. The nearest cloud seems to be located between 70 and 130 pc with A up to 0.7 and the next larger cloud between 250 and 500 pc with A up to 1.4. Based on constraints obtained form the modeling in the next section we find that DraChi is unlikely to be located in the first cloud because the inferred star luminosity is too low for an A star. It is also unlikely to be located beyond 350 pc because of the apparent size of the disk and large absorption values past this distance. We adopt a distance of 300 pc for our modeling and in the discussion section we present more arguments for choosing this distance.
5 Radiation Transfer Modeling
To investigate very approximate disk parameters based on the various data at hand we use the three-dimensional dust continuum Monte-Carlo radiative transfer code HOCHUNK3D666https://gemelli.colorado.edu/ bwhitney/codes/codes.html (Whitney et al., 2013). The basic structure of a typical flared disk can be described by the height of the disk, the surface density, and the density structure, respectively:
where is the disk scale height exponent, is the disk radial density exponent, and and are the scale height and surface density of the disk at the pivot radius, .
We first use the optical spectrum and the PS1 images to impose some constraints on the parameters. The optical spectrum is consistent with a star temperature in the range 6500 - 8500 K. The PS1 images constrain the inclination angle to 80 - 84 and the apparent size of the disk to 11.
Finally, based on the distance of 300 pc estimated in Sect. 4, we can estimate the width of the disk and the scale height constrained by the PS1 images. We find a scale height of 0.05 - 0.1 at a radius of 500 AU. The dust destruction radius as defined in Robitaille et al. (2006) is given by , with the dust sublimation temperature . For approximate values of and , or 0.25 AU, which we set as the inner radius of the disk.
Miotello et al. (2014) describe Class I objects as disks with partially dissipating envelopes that contain large grains embedded deep within the inner regions of the envelopes. Such an envelope is suggested by the faint “fangs” seen in the PS1 images. Our model for DraChi includes a disk with large grains settled closer to the mid-plane as well as the usual small grain disk plus an Ulrich envelope (Ulrich, 1976).
HOCHUNK3D allows for two coplanar disks with their own parameters for the structure and dust composition. One disk contains large dust grains with a small scale height, and the other small dust grains with a larger scale height. Both disks begin at the sublimation radius of 0.25 AU. The small grain disk ends at 1650 AU, whereas the large grain disk ends at 500 AU which best fits the near-IR. For the large grains, we used Model 1 of Wood et al. (2001), which combines a power-law with exponential cutoff for components of amorphous carbon and astronomical silicates with solar abundance constraints for both, and a max radius of a = 1000 m. For the small grains, we used the model of Kim et al. (1994), which is a mixture of silicates and graphite where the size distribution is determined by maximum-entropy fitting of the interstellar extinction curve. For the silicates, a = 0.16 m and for the graphite, a = 0.28 m. Both disks also include polycyclic aromatic hydrocarbons and the very small grains (PAHs/VSGs) dust model from Draine & Li (2007). The Ulrich envelope has the same composition as the small grains disk, has an ambient density at the inner sublimation radius of g/s and extends to 3000 AU.
Exhaustive radiation transfer modelling is beyond the scope of this letter; the intention here is instead to estimate a plausible set of coarse disk parameters that explain the data reasonably well. Given the constraints imposed by the PS1 image and the A9 spectral type, we explored a range of parameters to find a model that best fits the SED and especially the far-IR as it is not affected by extinction. The small and large grains scale heights are 0.06 and 0.01 at 500 AU, respectively. The scale height exponent is 1.15 for the small grain disk and 1.4 for the large grains. The radial density for both disks was adjusted to best fit both near-IR and far-IR data, with = 2.2 for both disks. The dust-to-gas ratio was taken at the standard fraction of 0.01.
The best model is presented in Figures 1 and 4 which show the model image convolved with the PS1 PSFs, and the model SED, respectively. We emphasise that this model is not unique, but it provides a good match to the available data. However, there are significant discrepancies in the near and mid-IR range which we discuss below. It is also remarkable how well the envelope included in the model resembles the faint “fangs” seen in the PS1 image in the northern part of the disk (Figure 1). The stellar effective temperature is 8000 K with a luminosity 11.46 L. The total disk mass is 0.2 M, with 20% of it in the large grains. The inclination angle is 82. A line-of-sight extinction was added to the model.
6 Discussion
6.1 DraChi: the largest known protoplanetary disk
DraChi is a newly-identified edge-on protoplanetary disk and in angular extent is the largest such object now known. This identification is based on PS1 images, and the results of radiation transfer modeling are also consistent with a late A-type star surrounded by a protoplanetary disk seen almost edge-on, and by an almost dissipated envelope.
DraChi is very similar to the well known edge-on protoplanetary disk GoHam, they both likely have HAeBe central stars, have a large apparent size and are not associated with known star clusters. One significant difference is the presence of the faint “fangs” in the PS1 images of DraChi, which we interpret as a dissipating envelope. DraChi is about 50% larger than GoHam in angular extent. We note that GoHam presents evidence for a possible protoplanet (Berné et al., 2015), showing that planet formation might be detected in such edge-on disks.
6.2 The SED
Modelling of the SED shown in Figure 4 presents a good overall match to observations, but discrepancies occur in the near- and mid-IR. We included PAH/VSG emisison in an attempt to reproduce the SED shape following Wood et al. (2008), but there are still significant discrepancies in mid-infrared around 10-20 m. We note that the photometric data is not simultaneous and such disks are known to be variable. The central star luminosity and temperature match a PMS star of type late A.
6.3 The “Fangs”
The radiative transfer model also reproduces well the PS1 image (Figure 1), including the faint “fangs” extending 18 in the northern part of the disk. The “fangs” are very similar to the “edge” of the envelope created by the shadow of the disk on the illuminated envelope. The southern part of this envelope is likely obscured in the PS1 images but it might be detected with more sensitive observations.
Another possibility is that the “fangs” are filaments associated with the disk itself. Smaller such filaments can be seen on the left side of the disk but these are located close to the central part of the disk, suggesting a possible outflow. If indeed the envelope is still present, this suggests a young disk at the end of the Class I phase. Another significant difference is the pronounced flaring seen in both lobes but most clearly in the southern part of the right lobe. The flaring also makes the left lobe curve outward, in contrast to the model image. Increasing the flaring in our model could not reproduce these features. Such flaring was seen in a similar HAeBe edge-on disk, PDS 144N (Perrin et al., 2006), which was also modelled with an envelope.
6.4 The Distance to DraChi
Estimating the distance to edge-on disks is hard if they are not associated with known star-forming regions, as demonstrated by GoHam with estimates between 250 and 500 pc, or PDS 144N (Perrin et al., 2006). The same is true for DraChi, and here we present arguments for preferring the adopted distance of 300 pc as opposed to a distance smaller than 150 pc, corresponding to the closest cloud found in Section 4, or larger distances of 350 - 500 pc.
Firstly, for a distance of 300 pc the disk radius is 1650 AU, which is already at the higher end of the range of radii for known disks. For larger distances of 350 - 500 pc the disk radius should be 1925 - 2750 pc which is unlikely. Secondly, the extinction estimated from Gaia for distances corresponding to the nearest cloud in Figure 3 are smaller than 0.7 A, and those for the next cloud go up to 1.4 A. Both our estimates based on the star spectrum and the radiation transfer modeling are close to 0.9 A which is the extinction corresponding to 300 pc in Figure 3, indicating that DraChi is likely in the second cloud but at the front part of it. Finally, if DraChi is located in the nearest cloud the luminosity of the star should be too small for an A star. For 300 pc the modeling suggests a luminosity of 11.5 L, but for distances less than 150 pc the star luminosity should be less than 5.0 L which is unlikely for a HAeBe star (Montesinos et al., 2009). We should note, however, that our estimated stellar type is based on scattered light alone as the star is not seen directly and that the observed luminosity in edge-on disks depends on multiple factors such as flaring parameter, disk mass, etc. Finally, a similar argument is based on the mass of the disk estimated from the radiation transfer modeling. The disk mass is given mostly by the far-infrared peak seen on the SED so it is not affected much by extinction but is also determined by the disk geometry such as flaring. Our model implies a disk mass of . If the distance is larger, the disk mass estimate increases to over for a distance of 500 pc, which again becomes difficult to explain in the context of known protoplanetary disk parameters.
For all these reasons we prefer the distance of 300 pc, and consider a distance smaller than 150 pc or larger than 350 pc unlikely.
@@ 7 Conclusion
We have identified a new, very large protoplanetary disk from PS1 images and confirmed the consistency of available data with a disk interpretation through radiative transfer and SED modelling. The host star was estimated to be an A9 HAeBe star based on an archival optical spectrum. We use the PS1 image to constrain the radiation transfer model parameters and the SED to find the best model. While the data is insufficient to perform detailed modeling, our model matches the observed SED well, although discrepancies are present in the mid-IR. The model disk has a mass of 0.2 , an inclination angle of 82, and its radius is 1650 AU at the adopted distance of 300 pc estimated from Gaia stellar extinctions. The low-density envelope included in the model has a 3000 AU radius and reproduces well the faint “fangs” seen in the northern part of the disk. If this structure is indeed part of the envelope, it suggests a young system at the end of the Class I phase. The serendipitous discovery of DraChi implies that such edge-on disks not associated with known star-forming regions (GoHam is another case) are still waiting to be discovered.
@@
| 3.4 | 57.3 1.21 | Cutri et al. (2014) | WISE | 6.1 |
| 4.6 | 116 2.14 | Cutri et al. (2014) | WISE | 6.4 |
| 12 | 310 3.10 | Cutri et al. (2014) | WISE | 6.5 |
| 22 | 318 5.27 | Cutri et al. (2014) | WISE | 12.0 |

### v2

@@ 5 Distance
Since DraChi is not associated with any known star-forming region the distance cannot be precisely determined, similar to GoHam which has estimates ranging from 250 to 500 pc or PDS 144N with an even wider range of 140 to 2000 pc (Perrin et al., 2006). However, the galactic coordinates of DraChi (l = 113.4, b = 6.4 ) places it close to the Cepheus Flare Shell. This is thought to be an expanding supernova bubble and is home to a complex of molecular clouds and star-forming regions. It is located at l = 120 and b = 17, a distance of 300 pc and is 9.5 in radius. Recently, Gaia data has been used to measure the distances to some of the dark clouds and young stars in this region (Sharma et al., 2022; Szilágyi et al., 2021) with values ranging from 150 to 900 pc, with the majority of stars between 330 and 370 pc. These methods used parallaxes and proper motions of known young stars to identify different star formation groups. While such methods are beyond the scope of this letter, we used Gaia extinctions which are part of the DR3 release to look for molecular clouds in the direction of DraChi. Molecular clouds increase the extinction of the stars behind them, thus producing breakpoints in absorption along the line of sight, therefore clouds can be found by identifying the breakpoints in absorption (Yan et al., 2019).
We selected stars located within 3 distance from DraChi. The Gaia extinctions were converted to AV using the dustapprox tool666https://github.com/mfouesneau/dustapprox. The result is shown in Figure 4. The nearest cloud seems to be located between 70 and 130 pc with AV up to 0.7 and the next larger cloud between 250 and 500 pc with AV up to 1.4. In the modeling section we adopted the distance of 300 pc which places DraChi at the front of the second cloud. We give below some qualitative arguments for preferring this distance but we emphasize this is not a precise measurement by any means. We expect more data or more sophisticated methods as mentioned above can improve this estimate. Some of these arguments are based on the modeling results so we used our best model to run simulations for distances of 100 and 500 pc while keeping the physical parameters the same to compare the results with those for the adopted distance of 300 pc.
First we compare the disk mass () and size (1650 AU) we obtained from the best model (at 300 pc) with the values of other known disks (Williams & Cieza, 2011; Andrews, 2020). Most known protoplanetary disks have masses and sizes 1000 AU so our estimates are already very high compared to other disks. Natta et al. (2000) investigated A-type PMS stars in particular, they show that for stars with masses the disks have masses . The same disk mass limit is measured by Boissier et al. (2011) for stars with masses . For a distance of 500 pc, the disk radius is 2750 AU and the modeling gives a disk mass . The mass ratio disk/star is in this case larger than 0.25. Secondly, the extinction estimated from Gaia for distances corresponding to the nearest cloud in Figure 4 are smaller than 0.7 AV, and those for the next cloud go up to 1.4 AV. Both our estimates based on the star spectrum and the radiative transfer modeling are 0.95 AV which is approximately the extinction that corresponds to 300 pc in Figure 4. The estimate from the spectrum includes both the intrinsic and the line-of-sight extinction so it is even less likely that the line-of-sight extinction is larger than 0.9 AV but we cannot rule out the a lower extinction which corresponds to the nearest cloud. We also note that the absorption estimated from the spectrum is based on the reddening and not the total extinction which is expected to be very high. The total absorbed luminosity estimated from the best model is about 2.5 L⊙ and the estimated luminosity of the star is 11.46 L⊙ with most of the extinction in the optical where the disk is optically thick. Both these arguments point against a distance as high as 500 pc.
The final argument is against a lower distance of about 100 pc, based on the star luminosity obtained from the modeling and the spectral type of the star obtained from the optical spectrum. For our 300 pc model we obtained a luminosity of 11.5 L⊙ but for a distance of 100 pc the star luminosity obtained from the model would be 5.2 L⊙, which is low for a HAeBe star (Montesinos et al., 2009). We point out that these estimates of the star luminosity are based on fitting the model to the SED and in particular to the mid and far-IR which are not affected much by extinction. However, there other factors that contribute to the luminosity estimates such as disk mass or flaring.
For all these reasons we prefer the distance of 300 pc but we emphasize again that this is not a precise estimate.
6 DraChi: the largest known protoplanetary disk
DraChi is one of only three known HAeBe edge-on protoplanetary disks. As already pointed out, it shares quite a few similarities with GoHam, including a large apparent size, but DraChi is about 50% larger. Located a distance of 250-300 pc, the star in GoHam is surrounded by a 0.3 disk with an inclination of (Wood et al., 2008; Bujarrabal et al., 2008). These values are quite close to our estimates for DraChi, and we followed the radiative transfer modeling of GoHam in Wood et al. (2008), including the dust properties. We note that GoHam presents evidence for a possible protoplanet (Berné et al., 2015; Teague et al., 2020), showing that planet formation might be detected in such edge-on disks.
One important difference from DraChi is the absence of an envelope as GoHam is thought to be a more evolved disk, while DraChi appears to still be surrounded by a faint and probably dissipating envelope if our interpretation of the “fangs” is correct. Similar filaments or “rays” are present in HV Tau C, extending out to almost double the size of the disk. They are closer to the mid-plane and straighter compared to our “fangs”. Stapelfeldt et al. (2003) and Duchêne et al. (2010) give different interpretation for these filaments in HV Tau C, as part of the disk or the envelope, which is quite bright in this edge-on system. For DraChi we prefer the second option, the fangs are quite similar to the “edges” of the envelope in Figure 1. In addition, extended emission can be seen from the fainter lobe and additional faint filaments in the northern part of this lobe, extending perpendicular to the disk. Another bright envelope is present in the Class I disk called the Butterfly (Villenave et al., 2024), in this case the envelope dominates the scattered light emission and shows clear filaments extending out to large distances.
The third known HAeBe edge-on protoplanetary disk is PDS 144N wth an inclination of 83. This system has no conspicuous envelope but shows pronounced flaring at the edges of the disk, similar to what we see at the lower edge of the bright lobe of Drachi (Figure 1). Perrin et al. (2006) was able to simulate these “wings” with a model that included an envelope and cavity and a dust model which included PAHs.
Evidence for grain settling was found in many edge-on disks such as HH 30 IRS (Wood et al., 2002) and Oph163131 (Villenave et al., 2022). It is unclear how long the grain settling takes place, but recent studies suggest most settling takes place between Class I and II (Villenave et al., 2024) where we think DraChi is currently. Villenave et al. (2020) showed that some edge-on disks show evidence that large grain settling takes place at scale heights significantly smaller than the small grains. Our radiative transfer model simulated grain settling by adding a disk with large grains at a scale height six times smaller than the small grains. Constraining the grain size is very important in modeling disks, but requires observations in millimeter (Monsch et al., 2024). For example, it is well known that grain size distributions and opacities have a strong impact on disk mass estimates. If large grains are included compared to ISM type dust the mass estimates are increased significantly as the gas is decoupled from the dust (Wood et al., 2002). One example is HK Tau B (Duchêne et al., 2003), where estimates of the disk mass were shown to vary by a factor of 5 depending on the grains used.
We detected a lateral asymmetry in DraChi, similar to many other edge-on disks. Villenave et al. (2024) found that 15 out of 20 edge-on disks in their sample show clear lateral asymmetries, including GoHam and PDS 144N, and they attribute this effect to a tilted inner region of the disk. Such a disk also changes the brightness of the lobes in edge-on disks and can even lead to an inversion of the brightest lobes at specific wavelengths as they detect in mid-infrared in the Butterfly disk. We also found a significant lateral asymmetry in DraChi (panel d of Figure 1) and therefore our estimate of the disk inclination in Drachi based on the brightness of the lobes could be biased.
@@ 7 Conclusion
We have identified a new, very large protoplanetary disk from PS1 images and confirmed the consistency of available data with a disk interpretation through radiative transfer and SED modelling. The host star was estimated to be an A9 HAeBe star based on an archival optical spectrum. We use the PS1 image to constrain the radiative transfer model parameters and the SED to find the best model. Our data is insufficient to constrain the grain sizes and spatial distribution and we therefore employed a model used successfully on other edge-on disks including GoHam. Our model matches the observed SED reasonably well, although larger discrepancies are present in the mid-IR. The model disk has a mass of 0.2 , an inclination angle of 82, and its radius is 1650 AU at the adopted distance of 300 pc estimated from Gaia stellar extinctions. The low-density envelope included in the model has a 3000 AU radius and reproduces well the faint “fangs” seen in the northern part of the disk. If this structure is indeed part of the envelope, it suggests a young system at the end of the Class I phase. We found a lateral asymmetry in the disk as detected in many edge-on disks, suggesting a tilted inner region. The serendipitous discovery of DraChi implies that such edge-on disks not associated with known star-forming regions (GoHam is another case) are still waiting to be discovered.
@@
| 3.4 | 57.3 1.21 | Cutri et al. (2014) | ALLWISE | 6.1 |
| 4.6 | 116 2.14 | Cutri et al. (2014) | ALLWISE | 6.4 |
| 12 | 310 3.10 | Cutri et al. (2014) | ALLWISE | 6.5 |
| 22 | 318 5.27 | Cutri et al. (2014) | ALLWISE | 12.0 |